OpenStudy (anonymous):

First, solve the equation for y. Second, find the value of y for the given x-value. 3x - 2y = -7, when x = 3

OpenStudy (anonymous):

y = (3x+7)/2 when x = 3, 2y = 9+7 y = 16/2 = 8

OpenStudy (anonymous):

Thank you.

OpenStudy (anonymous):

Welcome :)